# Hidden-Heavy Pentaquarks and Where to Find Them

### v1

@@ Abstract
We provide a simple explanation for the observed hidden-charm pentaquarks as bound states in Born-Oppenheimer potentials. We identify $P_{c\bar{c}}(4312)^{+}$, $P_{c\bar{c}}(4380)^{+}$, $P_{c\bar{c}}(4440)^{+}$, and $P_{c\bar{c}}(4457)^{+}$ as the heavy-quark spin states in a quartet of $c\bar{c}$ pentaquarks with $J^{P}$ quantum numbers $\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, $\frac{3}{2}^{-}$, and $\frac{5}{2}^{-}$. The quantum numbers of $P_{c\bar{c}}(4457)^{+}$ differ from all previous predictions. We identify $P_{c\bar{c}s}(4338)^{0}$ and $P_{c\bar{c}s}(4459)^{0}$ as spin states in a triplet of $c\bar{c}s$ pentaquarks with quantum numbers $\frac{1}{2}^{-}$, $\frac{1}{2}^{-}$, and $\frac{3}{2}^{-}$. We explain why $P_{c\bar{c}}(4380)^{+}$ is the only observed hidden-charm pentaquark with a large width.
@@
Introduction. The unexpected discoveries of dozens of exotic heavy hadrons since 2003 has provided a major challenge to our understanding of quantum chromodynamics (QCD) [1, 2, 3, 4, 5, 6, 7, 8]. Constituent models, such as molecular models with color-singlet hadron constituents and quark and diquark models with colored constituents, have failed to reveal the pattern of the exotic heavy hadrons. The direct assault on this problem using lattice QCD seems to be prohibitively difficult because of their many decay channels.
A promising approach within QCD is the Born-Oppenheimer (B-O) approximation, which exploits the large mass of a heavy (charm or bottom) quark compared to the energies of gluons and light quarks. The B-O approximation for QCD was introduced by Juge, Kuti, and Morningstar, who applied it to quarkonium hybrids in 1999 [9]. It separates the problem into two steps: (1) the calculation of B-O potentials, which are discrete energy levels of QCD in the presence of static color sources as functions of the separations of the sources, and (2) the solution of the Schrödinger equation for heavy quarks and antiquarks interacting through those potentials. This approach has been developed into an effective field theory called BOEFT that can be applied to all multi-heavy hadrons [10, 11, 12]. However, the development of BOEFT has not revealed the pattern of the exotic heavy hadrons.
A pattern for the exotic hidden-heavy hadrons, which contain a heavy quark and antiquark, was proposed in Ref. [13]. BOEFT implies that their B-O potentials at short distances are repulsive color-Coulomb potentials offset by the energy of an $\bm{8}$-hadron, which is a discrete energy level of QCD in the presence of a static color-octet source. Confinement and cluster decomposition imply that their B-O potentials at long distances approach thresholds for pairs of heavy hadrons. Such potentials can support bound states only if they cross below the heavy-hadron-pair threshold before approaching it, which may happen only for the potentials associated with the lowest-energy $\bm{8}$-hadrons. The pattern proposed in Ref. [13] is that the exotic hidden-heavy hadrons are bound states and resonances in such B-O potentials. In Ref. [13], this pattern was applied to nonstrange hidden-heavy tetraquarks and some of their properties were postdicted. The LHCb collaboration has discovered four nonstrange hidden-charm ($c\bar{c}$) pentaquarks [14, 15, 16] and two strange hidden-charm ($c\bar{c}s$) pentaquarks [17, 18]. In this paper, we apply the pattern in Ref. [13] to hidden-heavy pentaquarks and we predict some of their properties.
@@
A $\bm{\bar{3}}$-hadron with a single quark is a $\bm{\bar{3}}$-meson. A $\bm{3}$-hadron with two quarks is a $\bm{3}$-baryon. At large $r$, a pentaquark B-O potential approaches the threshold given by the sum of the energies of a $\bm{\bar{3}}$-meson and a $\bm{3}$-baryon. The color, spin, and flavor states of the quarks in a static hadron are in the fundamental representations of SU(3), SU(2), and SU(3), respectively. The lowest-energy $\bm{\bar{3}}$-mesons are $S$-wave states with even parity ($\pi=+$). They therefore form a $j^{\pi}=\frac{1}{2}^{+}$ SU(3) triplet consisting of $(I,S)=\bigl{(}\frac{1}{2},0)$ and $(0,-1)$. The state of a $\bm{3}$-baryon must be antisymmetric under exchange of the two quarks. The lowest-energy $\bm{3}$-baryons are symmetric in the quark positions and therefore have $\pi=+$. The color state of the two quarks must be the antisymmetric antitriplet, so their spin-flavor state must be symmetric. The lowest-energy $\bm{3}$-baryons therefore form two SU(3) multiplets: a $j^{\pi}=0^{+}$ antitriplet consisting of $(I,S)=(0,0)$ and $\bigl{(}\frac{1}{2},-1\bigr{)}$ and a $1^{+}$ sextet consisting of $(1,0)$, $\bigl{(}\frac{1}{2},-1\bigr{)}$, and $(0,-2)$.
@@
A heavy hadron contains a single heavy quark or antiquark and light QCD fields that can be approximated by a $\bm{3}$-hadron or $\bm{\bar{3}}$-hadron. The heavy-quark spin is conserved up to terms that scale like $1/m_{Q}$ with $m_{Q}$ the heavy-quark mass. Heavy hadrons form approximately degenerate heavy-quark spin symmetry (HQSS) multiplets labeled by the quantum numbers $j^{\pi}$ and $(I,S)$ of a static hadron. The members of the multiplets are labeled by spin and parity quantum numbers $J^{P}$. The HQSS multiplet is a singlet with $J=\frac{1}{2}$ if $j=0$ or a doublet with $J=j-\frac{1}{2},j+\frac{1}{2}$ if $j\geq\frac{1}{2}$. The parity is $P=+\pi$ or $-\pi$ for a hadron with a heavy quark or antiquark.
@@
The $\frac{1}{2}^{+}$ $\bm{\bar{3}}$-meson triplet corresponds to HQSS doublets of charm mesons with $J^{P}=0^{-},1^{-}$: $\bar{D}^{(\ast)}\equiv(\bar{D},\bar{D}^{\ast})$ and $\bar{D}_{s}^{(\ast)}\equiv(\bar{D}_{s}\bar{D}_{s}^{\ast})$ with $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$ and $(0,-1)$. The $0^{+}$ $\bm{3}$-baryon antitriplet corresponds to charm baryons with $J^{P}=\frac{1}{2}^{+}$: $\Lambda_{c}$ and $\Xi_{c}$ with $(I,S)=(0,0)$ and $\bigl{(}\frac{1}{2},-1)$. The $1^{+}$ $\bm{3}$-baryon sextet corresponds to HQSS doublets of charm baryons with $J^{P}=\frac{1}{2}^{+},\frac{3}{2}^{+}$: $\Sigma_{c}^{(\ast)}\equiv(\Sigma_{c},\Sigma_{c}^{\ast})$, $\Xi_{c}^{\prime(\ast)}\equiv(\Xi_{c}^{\prime},\Xi_{c}^{\prime\ast})$, and $\Omega_{c}^{(\ast)}\equiv(\Omega_{c},\Omega_{c}^{\ast})$ with $(I,S)=(1,0)$, $\bigl{(}\frac{1}{2},-1\bigr{)}$, and $(0,-2)$.
@@
An $\bm{8}$-hadron with three quarks is an $\bm{8}$-baryon. At small $r$, a pentaquark B-O potential approaches a repulsive color-Coulomb potential offset by the energy of an $\bm{8}$-baryon. The B-O potentials that are most likely to support bound states are those associated with the lowest-energy $\bm{8}$-baryons. The state of an $\bm{8}$-baryon must be completely antisymmetric under exchanges of the three quarks. The lowest-energy $\bm{8}$-baryons are expected to be symmetric in the quark positions and therefore have $\pi=+$. The color state of the three quarks must be a mixed-symmetry octet, so their spin-flavor state must have mixed symmetry. The lowest-energy $\bm{8}$-baryons are expected to form four SU(3) multiplets: a $j^{\pi}=\frac{1}{2}^{+}$ singlet with $(I,S)=(0,-1)$, a $\frac{1}{2}^{+}$ octet and a $\frac{3}{2}^{+}$ octet each consisting of $\bigl{(}\frac{1}{2},0\bigr{)}$, $(0,-1)$, $(1,-1)$, and $\bigl{(}\frac{1}{2},-2\bigr{)}$, and a $\frac{1}{2}^{+}$ decuplet consisting of $\bigl{(}\frac{3}{2},0\bigr{)}$, $(1,-1)$, $\bigl{(}\frac{1}{2},-2\bigr{)}$, and $(0,-3)$.
@@
Pentaquark B-O Potentials. The B-O potentials for hidden-heavy hadrons are labeled by quantum numbers for the symmetries of QCD with static triplet and antitriplet sources. For sources separated by a vector $\bm{r}$, the symmetries are rotations around $\bm{\hat{r}}$, a reflection $R$ through any plane containing $\bm{\hat{r}}$, and $CP$. The traditional B-O quantum numbers are $\Lambda_{\eta}^{\epsilon}$, where $\Lambda=\lvert\bm{J}\cdot\bm{\hat{r}}\rvert$ with $\bm{J}$ the angular-momentum vector, $\eta=g$ or $u$ for even or odd $CP$, and $\epsilon=+$ or -- for even or odd $R$. For pentaquark B-O potentials, states with $\eta=g,u$ are equal-energy superpositions with opposite baryon numbers and states with $\epsilon=+,-$ are equal-energy superpositions with opposite $\bm{J}\cdot\bm{\hat{r}}$. Therefore, the subscript $\eta$ and the superscript $\epsilon$ can both be omitted and the B-O quantum numbers reduce simply to the half-integer $\Lambda$. The B-O potentials are also labeled by flavor quantum numbers $(I,S)$. The $c\bar{c}$ pentaquarks were discovered through their decays into $J/\psi\,p$, which implies $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$. The $c\bar{c}s$ pentaquarks were discovered through their decays into $J/\psi\,\Lambda$, which implies $(I,S)=(0,-1)$.
@@
At large $r$, the B-O potentials form degenerate multiplets labeled by the quantum numbers $j_{\text{B}}^{\pi_{\text{B}}}$, $(I_{\text{B}},S_{\text{B}})$ of a $\bm{3}$-baryon and $j_{\text{M}}^{\pi_{\text{M}}}$, $(I_{\text{M}},S_{\text{M}})$ of a $\bm{\bar{3}}$-meson, which can be alternatively specified by the HQSS multiplets of heavy baryons and heavy mesons. Each multiplet of B-O potentials can be decomposed into sub-multiplets labeled by $j^{\pi}$ with $j=\lvert j_{\text{B}}-j_{\text{M}}\rvert,\dots,j_{\text{B}}+j_{\text{M}}$ and $\pi=\pi_{\text{B}}\pi_{\text{M}}$. We label these potentials by $(j^{\pi})\Lambda$ with $\Lambda=\frac{1}{2},\dots,j$. Their isospin is $I=\lvert I_{\text{B}}-I_{\text{M}}\rvert,\dots,I_{\text{B}}+I_{\text{M}}$ and their strangeness is $S=S_{\text{B}}+S_{\text{M}}$. Some of the B-O potentials with the lowest energy at large $r$ are listed in Table 1.
@@
Table 1: The lowest-energy pairs of a $\bm{3}$-baryon and a $\bm{\bar{3}}$-meson with $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$ and $(0,-1)$ and the associated B-O potentials $(j^{\pi})\Lambda$ at large $r$.
Static-hadron pair $I$ $S$ $(j^{\pi})\Lambda$
@@
$\Lambda_{c}\bar{D}_{s}^{(\ast)}$ 0 $-1\hphantom{-}$ $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$
$\Xi_{c}\bar{D}^{(\ast)}$ 0 $-1\hphantom{-}$ $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$
$\Xi_{c}^{\prime(\ast)}\bar{D}^{(\ast)}$ 0 $-1\hphantom{-}$ $\bigl{(}\frac{3}{2}^{+}\bigr{)}\frac{1}{2}$, $\bigl{(}\frac{3}{2}^{+}\bigr{)}\frac{3}{2}$, $\bigl{(}\frac{1}{2}^{+\prime\prime}\bigr{)}\frac{1}{2}$
At small $r$, the B-O potentials form degenerate multiplets labeled by the quantum numbers $j^{\pi}$ and $(I,S)$ of an $\bm{8}$-baryon. We label these potentials by $(j^{\pi})\Lambda$ with $\Lambda=\frac{1}{2},\dots,j$. Their flavor is that of the $\bm{8}$-baryon. Some of the B-O potentials with the lowest energy at small $r$ are listed in Table 2.
Table 2: The lowest-energy $\bm{8}$-baryons with $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$ and $(0,-1)$ and the associated B-O potentials $(j^{\pi})\Lambda$ at small $r$.
$\bm{8}$-baryon $I$ $S$ $(j^{\pi})\Lambda$
@@
$\frac{1}{2}^{+}$ 0 $-1\hphantom{-}$ $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$
$\hphantom{{}^{\prime}}\frac{1}{2}^{+\prime}$ 0 $-1\hphantom{-}$ $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$
$\frac{3}{2}^{+}$ 0 $-1\hphantom{-}$ $\bigl{(}\frac{3}{2}^{+}\bigr{)}\frac{1}{2}$, $\bigl{(}\frac{3}{2}^{+}\bigr{)}\frac{3}{2}$
@@
Since the spectrum of QCD with two static color sources must be a smooth function of $r$, the B-O potentials associated with a $\bm{3}$-baryon and $\bm{\bar{3}}$-meson at large $r$ must connect smoothly to the B-O potentials associated with an $\bm{8}$-baryon at small $r$. The $(j^{\pi})\Lambda$ potential with the $n$th lowest energy at large $r$ must connect to the B-O potential with the same $\Lambda$ and $(I,S)$ and the $n$th lowest energy at small $r$. The simplest possibility is that the quantum numbers $j^{\pi}$ at small $r$ coincide with those at large $r$. Under this assumption, we proceed to connect the B-O potentials in Table 1 to those in Table 2 and identify the potentials that may support bound states.
@@
First, we examine B-O potentials with $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$. The $\Lambda_{c}\bar{D}^{(\ast)}$ potential $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$ must connect to the potential for the $\frac{1}{2}^{+}$ $\bm{8}$-baryon. The two $\Sigma_{c}^{(\ast)}\bar{D}^{(\ast)}$ potentials $\bigl{(}\frac{3}{2}^{+}\bigr{)}\Lambda$ must connect to the two potentials for the $\frac{3}{2}^{+}$ $\bm{8}$-baryon. The $\Sigma_{c}^{(\ast)}\bar{D}^{(\ast)}$ potential $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ must connect to the potential for an orbitally-excited $\bm{8}$-baryon, since Table 2 does not have a second $\bm{8}$-baryon with $j^{\pi}=\frac{1}{2}^{+}$ and $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$. The lack of evidence for $c\bar{c}$ pentaquarks near the $\Lambda_{c}\bar{D}$ and $\Lambda_{c}\bar{D}^{\ast}$ thresholds suggests that the $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$ potential does not support bound states. Each of the four observed $c\bar{c}$ pentaquarks is near one of the $\Sigma_{c}\bar{D}$, $\Sigma_{c}^{\ast}\bar{D}$, and $\Sigma_{c}\bar{D}^{\ast}$ thresholds. This suggests that either the $\bigl{(}\frac{3}{2}^{+}\bigr{)}\Lambda$ potentials or the $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ potential or both support bound states. It is implausible that the $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ potential supports bound states, since it connects to the potential of an orbitally-excited $\bm{8}$-baryon. In this case, only the coupled $\bigl{(}\frac{3}{2}^{+}\bigr{)}\Lambda$ potentials would support bound states.

### v3

@@ Abstract
We provide a simple explanation for the observed hidden-charm pentaquarks as bound states in Born-Oppenheimer potentials. We identify $P_{c\bar{c}}(4312)^{+}$, $P_{c\bar{c}}(4440)^{+}$, and $P_{c\bar{c}}(4457)^{+}$ as heavy-quark spin states in a quartet of $c\bar{c}$ pentaquarks with $J^{P}$ quantum numbers $\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, and $\frac{5}{2}^{-}$. The quantum numbers of $P_{c\bar{c}}(4457)^{+}$ differ from most previous predictions. We also predict a fourth $c\bar{c}$ pentaquark with quantum numbers $\frac{3}{2}^{-}$ near the $\Sigma_{c}^{\ast}\bar{D}$ threshold. We identify $P_{c\bar{c}s}(4338)^{0}$ and $P_{c\bar{c}s}(4459)^{0}$ as heavy-quark spin states in a triplet of $c\bar{c}s$ pentaquarks with quantum numbers $\frac{1}{2}^{-}$ and either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$. We also predict a third $c\bar{c}s$ pentaquark with quantum numbers either $\frac{3}{2}^{-}$ or $\frac{1}{2}^{-}$ near the $\Xi_{c}\bar{D}^{\ast}$ threshold. We explain why the observed hidden-charm pentaquarks have narrow widths.
@@
Introduction. The unexpected discoveries of dozens of exotic heavy hadrons since 2003 has provided a major challenge to our understanding of quantum chromodynamics (QCD) [1, 2, 3, 4, 5, 6, 7, 8, 9]. Constituent models, such as molecular models with color-singlet hadron constituents and quark and diquark models with colored constituents, have failed to reveal the pattern of the exotic heavy hadrons. The direct assault on this problem using lattice QCD seems to be prohibitively difficult because of their many decay channels.
A promising approach within QCD is the Born-Oppenheimer (B-O) approximation, which exploits the large mass of a heavy (charm or bottom) quark compared to the energies of gluons and light quarks. The B-O approximation for QCD was introduced by Juge, Kuti, and Morningstar, who applied it to quarkonium hybrids in 1999 [10]. It separates the problem into two steps: (1) the calculation of B-O potentials, which are discrete energy levels of QCD in the presence of static color sources as functions of the separation of the sources $r$, and (2) the solution of the Schrödinger equation for heavy quarks and antiquarks interacting through those potentials. This approach has been developed into an effective field theory called BOEFT that can be applied to all multi-heavy hadrons [11, 12, 13]. However, the development of BOEFT has not revealed the pattern of the exotic heavy hadrons.
A pattern for the exotic hidden-heavy hadrons, which contain a heavy quark and antiquark, was proposed in Ref. [14]. BOEFT implies that their B-O potentials at short distances are repulsive color-Coulomb potentials offset by the energy of an $\bm{8}$-hadron, which is a discrete energy level of QCD in the presence of a static color-octet source. Confinement implies that their B-O potentials at long distances approach thresholds for pairs of heavy hadrons. Such potentials can support bound states only if they cross below the heavy-hadron-pair threshold before approaching it; see Figure 2. The pattern proposed in Ref. [14] is that the exotic hidden-heavy hadrons are bound states and resonances in such B-O potentials. In Ref. [14], this pattern was applied to nonstrange hidden-heavy tetraquarks and some of their properties were postdicted. The LHCb collaboration has discovered three nonstrange hidden-charm ($c\bar{c}$) pentaquarks [15, 16, 17] and two strange hidden-charm ($c\bar{c}s$) pentaquarks [18, 19]. [^1] We disregard the broad $c\bar{c}$ pentaquark state $P_{c\bar{c}}(4380)^{+}$ in the Particle Listings of the Particle Data Group for the reasons indicated in the PDG Review on Pentaquarks [20]. In this paper, we apply the pattern in Ref. [14] to hidden-heavy pentaquarks and we predict some of their properties.
@@
Figure 2: A B-O potential for exotic hidden-heavy hadrons (thick curve) may support bound states (thin horizontal line) only if it crosses below the heavy-hadron-pair threshold before approaching it.
@@
A $\bm{\bar{3}}$-hadron with a single light quark is a $\bm{\bar{3}}$-meson. A $\bm{3}$-hadron with two light quarks is a $\bm{3}$-baryon. At large $r$, a pentaquark B-O potential approaches the threshold given by the sum of the energies of a $\bm{\bar{3}}$-meson and a $\bm{3}$-baryon. The color, spin, and flavor states of the quarks in a static hadron are in the fundamental representations of SU(3), SU(2), and SU(3), respectively. The ground-state $\bm{\bar{3}}$-mesons are $S$-wave states with even parity ($\pi=+$). They therefore form a $j^{\pi}=\frac{1}{2}^{+}$ SU(3) triplet consisting of $(I,S)=\bigl(\frac{1}{2},0\bigr)$ and (0,1)(0,-1). The state of a $\bm{3}$-baryon must be antisymmetric under exchange of the two quarks. The ground-state $\bm{3}$-baryons are symmetric in the quark positions and therefore have $\pi=+$. The color state of the two quarks must be the antisymmetric antitriplet, so their spin-flavor state must be symmetric. The ground-state $\bm{3}$-baryons therefore form two SU(3) multiplets: a $j^{\pi}=0^{+}$ antitriplet consisting of $(I,S)=(0,0)$ and $\bigl(\frac{1}{2},-1\bigr)$ and a $1^{+}$ sextet consisting of $(1,0)$, $\bigl(\frac{1}{2},-1\bigr)$, and (0,2)(0,-2).
@@
A heavy hadron contains a single heavy quark or antiquark and light QCD fields that can be approximated by a $\bm{3}$-hadron or $\bm{\bar{3}}$-hadron. The heavy-quark spin is conserved up to terms that scale like $1/m_{Q}$ with $m_{Q}$ the heavy-quark mass. Heavy hadrons form approximately degenerate heavy-quark spin-symmetry (HQSS) multiplets labeled by the quantum numbers $j^{\pi}$ and $(I,S)$ of a static hadron. The members of the multiplets are labeled by spin and parity quantum numbers $J^{P}$. The HQSS multiplet is a singlet with $J=\frac{1}{2}$ if $j=0$ or a doublet with $J=j-\frac{1}{2},j+\frac{1}{2}$ if $j\geq\frac{1}{2}$. The parity is $P=+\pi$ or $\pi$ for a hadron with a heavy quark or antiquark.
@@
The $\frac{1}{2}^{+}$ $\bm{\bar{3}}$-meson triplet corresponds to HQSS doublets of charm mesons with $J^{P}=0^{-},1^{-}$: $\bar{D}^{(\ast)}\equiv(\bar{D},\bar{D}^{\ast})$ and $\bar{D}_{s}^{(\ast)}\equiv(\bar{D}_{s}\bar{D}_{s}^{\ast})$ with $(I,S)=\bigl(\frac{1}{2},0\bigr)$ and (0,1)(0,-1). The $0^{+}$ $\bm{3}$-baryon antitriplet corresponds to charm baryons with $J^{P}=\frac{1}{2}^{+}$: $\Lambda_{c}$ and $\Xi_{c}$ with $(I,S)=(0,0)$ and $\bigl(\frac{1}{2},-1\bigr)$. The $1^{+}$ $\bm{3}$-baryon sextet corresponds to HQSS doublets of charm baryons with $J^{P}=\frac{1}{2}^{+},\frac{3}{2}^{+}$: $\Sigma_{c}^{(\ast)}\equiv(\Sigma_{c},\Sigma_{c}^{\ast})$, $\Xi_{c}^{\prime(\ast)}\equiv(\Xi_{c}^{\prime},\Xi_{c}^{\prime\ast})$, and $\Omega_{c}^{(\ast)}\equiv(\Omega_{c},\Omega_{c}^{\ast})$ with $(I,S)=(1,0)$, $\bigl(\frac{1}{2},-1\bigr)$, and (0,2)(0,-2).
@@
An $\bm{8}$-hadron with three light quarks is an $\bm{8}$-baryon. At small $r$, a pentaquark B-O potential approaches a repulsive color-Coulomb potential offset by the energy of an $\bm{8}$-baryon. The B-O potentials that are most likely to support bound states are those associated with the ground-state $\bm{8}$-baryons. The state of an $\bm{8}$-baryon must be completely antisymmetric under exchanges of the three quarks. The ground-state $\bm{8}$-baryons are expected to be symmetric in the quark positions and therefore have $\pi=+$. The color state of the three quarks must be a mixed-symmetry octet, so their spin-flavor state must have mixed symmetry. The ground-state $\bm{8}$-baryons are expected to form four SU(3) multiplets: a $j^{\pi}=\frac{1}{2}^{+}$ singlet with (I,S)=(0,1)(I,S)=(0,-1), a $\frac{1}{2}^{+}$ octet and a $\frac{3}{2}^{+}$ octet each consisting of $\bigl(\frac{1}{2},0\bigr)$, (0,1)(0,-1), (1,1)(1,-1), and $\bigl(\frac{1}{2},-2\bigr)$, and a $\frac{1}{2}^{+}$ decuplet consisting of $\bigl(\frac{3}{2},0\bigr)$, (1,1)(1,-1), $\bigl(\frac{1}{2},-2\bigr)$, and (0,3)(0,-3).
@@
Pentaquark B-O Potentials. The B-O potentials for hidden-heavy hadrons are labeled by quantum numbers for the symmetries of QCD with static triplet and antitriplet sources. For sources separated by a vector $\bm{r}$, the symmetries are rotations around $\bm{\hat{r}}$, a reflection $R$ through any plane containing $\bm{\hat{r}}$, and $CP$. The traditional B-O quantum numbers are $\Lambda_{\eta}^{\epsilon}$, where $\Lambda=\lvert\bm{J}\cdot\bm{\hat{r}}\rvert$ with $\bm{J}$ the angular-momentum vector, $\eta=g$ or $u$ for even or odd $CP$, and $\epsilon=+$ or - for even or odd $R$. For pentaquark B-O potentials, states with $\eta=g,u$ are equal-energy superpositions with opposite baryon numbers and states with $\epsilon=+,$- are equal-energy superpositions with opposite $\bm{J}\cdot\bm{\hat{r}}$. Therefore, the subscript $\eta$ and the superscript $\epsilon$ can both be omitted and the B-O quantum numbers reduce simply to the half-integer $\Lambda$. The B-O potentials are also labeled by flavor quantum numbers $(I,S)$. The $c\bar{c}$ pentaquarks were discovered through their decays into $J/\psi\,p$, which implies $(I,S)=\bigl(\frac{1}{2},0\bigr)$. The $c\bar{c}s$ pentaquarks were discovered through their decays into $J/\psi\,\Lambda$, which implies (I,S)=(0,1)(I,S)=(0,-1).
@@
At large $r$, the B-O potentials form degenerate multiplets labeled by the quantum numbers $j_{\text{B}}^{\pi_{\text{B}}}$, $(I_{\text{B}},S_{\text{B}})$ of a $\bm{3}$-baryon and $j_{\text{M}}^{\pi_{\text{M}}}$, $(I_{\text{M}},S_{\text{M}})$ of a $\bm{\bar{3}}$-meson, which can be alternatively specified by the HQSS multiplets of heavy baryons and heavy mesons. Each multiplet of B-O potentials can be decomposed into sub-multiplets labeled by $j^{\pi}$ with $j=\lvert j_{\text{B}}-j_{\text{M}}\rvert,\dots,j_{\text{B}}+j_{\text{M}}$ and $\pi=\pi_{\text{B}}\pi_{\text{M}}$. We label these potentials by $(j^{\pi})\Lambda$ with $\Lambda=\frac{1}{2},\dots,j$. Their isospin is $I=\lvert I_{\text{B}}-I_{\text{M}}\rvert,\dots,I_{\text{B}}+I_{\text{M}}$ and their strangeness is $S=S_{\text{B}}+S_{\text{M}}$. The B-O potentials with $(I,S)=\bigl(\frac{1}{2},0\bigr)$ and (0,1)(0,-1) that have the lowest energy at large $r$ are listed in Table 1.
@@
Table 1: The pairs of a ground-state $\bm{3}$-baryon and a ground-state $\bm{\bar{3}}$-meson with $(I,S)=\bigl(\frac{1}{2},0\bigr)$ and (0,1)(0,-1) and the associated B-O potentials $(j^{\pi})\Lambda$ at large $r$.
Static-hadron pair II $S$ $(j^{\pi})\Lambda$
@@
$\Lambda_{c}\bar{D}_{s}^{(\ast)}$ 0 $1\hphantom{-}$ $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$
$\Xi_{c}\bar{D}^{(\ast)}$ 0 $1\hphantom{-}$ $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$
$\Xi_{c}^{\prime(\ast)}\bar{D}^{(\ast)}$ 0 $1\hphantom{-}$ $\bigl(\frac{3}{2}^{+}\bigr)\frac{1}{2}$, $\bigl(\frac{3}{2}^{+}\bigr)\frac{3}{2}$, $\bigl(\frac{1}{2}^{+\prime\prime}\bigr)\frac{1}{2}$
At small $r$, the B-O potentials form degenerate multiplets labeled by the quantum numbers $j^{\pi}$ and $(I,S)$ of an $\bm{8}$-baryon. We label these potentials by $(j^{\pi})\Lambda$ with $\Lambda=\frac{1}{2},\dots,j$. Their flavor is that of the $\bm{8}$-baryon. The B-O potentials with $(I,S)=\bigl(\frac{1}{2},0\bigr)$ and (0,1)(0,-1) that have the lowest energy at small $r$ are listed in Table 2.
Table 2: The ground-state $\bm{8}$-baryons with $(I,S)=\bigl(\frac{1}{2},0\bigr)$ and (0,1)(0,-1) and the associated B-O potentials $(j^{\pi})\Lambda$ at small $r$.
$\bm{8}$-baryon II $S$ $(j^{\pi})\Lambda$
@@
$\frac{1}{2}^{+}$ 0 $1\hphantom{-}$ $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$
$\hphantom{{}^{\prime}}\frac{1}{2}^{+\prime}$ 0 $1\hphantom{-}$ $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$
$\frac{3}{2}^{+}$ 0 $1\hphantom{-}$ $\bigl(\frac{3}{2}^{+}\bigr)\frac{1}{2}$, $\bigl(\frac{3}{2}^{+}\bigr)\frac{3}{2}$
@@
Since the spectrum of QCD with two static color sources must be a smooth function of $r$, the B-O potentials associated with a $\bm{3}$-baryon and $\bm{\bar{3}}$-meson at large $r$ must connect smoothly to the B-O potentials associated with an $\bm{8}$-baryon at small $r$. The $(j^{\pi})\Lambda$ potential with the nth lowest energy at large $r$ must connect to the B-O potential with the same $\Lambda$ and $(I,S)$ and the nth lowest energy at small $r$. The simplest possibility is that the quantum numbers $j^{\pi}$ at small $r$ coincide with those at large $r$. Under this assumption, we proceed to connect the B-O potentials in Table 1 to those in Table 2. We then use these connections together with experimental evidence to infer the potentials that may support bound states.
@@
First, we examine B-O potentials with $(I,S)=\bigl(\frac{1}{2},0\bigr)$. The $\Lambda_{c}\bar{D}^{(\ast)}$ potential $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$ must connect to the potential for the $\frac{1}{2}^{+}$ $\bm{8}$-baryon. The two $\Sigma_{c}^{(\ast)}\bar{D}^{(\ast)}$ potentials $\bigl(\frac{3}{2}^{+}\bigr)\Lambda$ must connect to the two potentials for the $\frac{3}{2}^{+}$ $\bm{8}$-baryon. The $\Sigma_{c}^{(\ast)}\bar{D}^{(\ast)}$ potential $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ must connect to the potential for an orbitally excited $\bm{8}$-baryon, since Table 2 does not have a second $\bm{8}$-baryon with $j^{\pi}=\frac{1}{2}^{+}$ and $(I,S)=\bigl(\frac{1}{2},0\bigr)$. The lack of evidence for $c\bar{c}$ pentaquarks near the $\Lambda_{c}\bar{D}$ and $\Lambda_{c}\bar{D}^{\ast}$ thresholds suggests that the $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$ potential does not support bound states. One of the three observed $c\bar{c}$ pentaquarks is near the $\Sigma_{c}\bar{D}$ threshold and the other two are near the $\Sigma_{c}\bar{D}^{\ast}$ threshold. This suggests that either the $\bigl(\frac{3}{2}^{+}\bigr)\Lambda$ potentials or the $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ potential or both support bound states. It is implausible that the $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ potential supports bound states, since it connects to the potential of an orbitally-excited $\bm{8}$-baryon. In this case, only the $\bigl(\frac{3}{2}^{+}\bigr)\Lambda$ potentials would support bound states.